When 25.0 mL of 0.400 M nitric acid and 25.0 mL of 0.200 M barium hydroxide are combined, the pH of the resulting solution will be : .....

A. Less than 7
B. Equal to 7
C. Greater than 7
Answer #1

Millimoles of HNO3 = 25×0.400 = 10

Millimoles of H+ = 10

Millimoles of barium hydroxide = 25×0.200 = 5

Millimoles of OH- ion = 2×5 = 10.

Thus both H+ and OH- neutralize each other and pH is 7

Hence, answer is B.